# Mojito: Motion Trajectory and Intensity Control for Video Generation

### v1

@@ 4.3 Model Backbone
The backbone of Mojito is based on a Conv-Spatial-Temporal Transformer architecture. The model structure consists of input blocks, a mid-block, and output blocks, each containing convolutional layers, temporal transformers, and spatial transformers. This design enables handling motion control in both the temporal and spatial dimensions.
The Temporal Transformer processes video data along the time axis, capturing dependencies between frames by reshaping the input into temporal sequences. This block includes relative position encoding and optional causal masking for unidirectional attention, making it effective for capturing motion patterns across frames. Group normalization (GroupNorm) ensures stability in representation learning.
The Spatial Transformer captures spatial relationships within each frame. Using linear or convolutional projections, this block reshapes and processes spatial information to learn localized patterns, which is essential for high-resolution video generation. Cross-attention layers enable alignment with external context when available, further refining spatial coherence in the generated content. The Directional Motion Control module works on this module.
Together, the temporal and spatial transformers enable Mojito to control motion directions and intensities. The details of the model architecture is given in the Supplementary Materials.
@@ 5.1 Datasets
We use Panda70M [8], InternVID [42], Mira [21] to train the model backbone and the Motion Intensity Modulator module. For Panda-70M, from which we collected a high quality training-10M subset. We processed another 1,000 videos for evaluation purpose for the motion intensity modulaton. For motion direction guidance evaluation, we use the LaSOT [9] dataset which contains prompt-bbox-video annotations; For InternVID and Mira, we processed and used the whole dataset for training. The details are given in the Supplementary materials.
Figure 3: Qualitative comparison of directional control with baselines.  Mojito achieves accurate control over object motion by specifying objects and precise locations without requiring additional training. The red bounding boxes, serving as inputs to Mojito, guide the generated balloon to follow the specified trajectory.
@@ 5.2 Evaluation Metrics
For quantitative evaluation, we employ commonly used metrics: Fréchet Video Distance (FVD) [39] and CLIPSIM [44], which calculates the average similarity across all generated frames in relation to the input caption. Specifically, we compute the FVD for a set of 300 videos, each consisting of 16 frames.
To quantify the efficacy of direction guidance, we use the off-the-shelf OWL-ViT-large open-vocabulary object detector [25] to detect the bounding box annotations of generated videos. We compute the Intersection-over-Union [33] of the detected bboxes and the input bboxes on videos. We also compute the Centroid Distance (CD) as the distance between the centroid of generated objects and input bboxes, normalized to 1. Finally, we report the average precision@50% (AP50) of the detected and input bboxes; For motion intensity control, we compute the difference between the detected average optical flow in the generated video and the target motion intensity, normalized to 1.
@@ 5.3 Implementation Details
Model backbone.
For the Variational Autoencoder (VAE) framework, the encoder processes input frames at a resolution of 320$\times$512, progressively downsampling through a series of residual blocks and convolutional layers with channel multipliers of $(1,2,4,4)$. This results in a latent representation with a spatial resolution of $40\times 64$. We use CLIP (ViT-H-14) [30] as the text encoder, allowing for high-quality text-to-video alignment. For video processing, spatial and temporal downsampling factors of 8 are applied in both dimensions. The attention matrices (such as $q$ and $kv$) within the spatial transformer are structured as $(t,b,c)$, where $t$ represents the temporal length. Trajectory adjustments are applied only within the mid-block and the initial layers of the upsampling branch in the denoising U-Net [35], which provides a balance between maintaining video quality and achieving precise motion control. Additional implementation details and parameter settings are provided in the Supplementary Materials.
Training.
We use the DeepSpeed [32] framework to train our model, utilizing memory and computation optimizations for distributed training. The model training is carried out using the FusedLamb [48] optimizer combined with the LambdaLRScheduler. We use a batch size of 4 and a base learning rate of $1.0\times 10^{-5}$. More details can be found in the Supplementary Materials.
@@ 5.4 Performance Evaluation
Table 1: Human evaluation between Mojito, OpenSora [52], OpenSora plan [23], and VideoCrafter2 [5]. It includes 2400 comparisons on 400 video pairs. Each video pair contains 5 ratings from different human annotators on Amazon Turk. The detailed setup is given in the Supplementary Materials.
Mojito Tie OpenSora [52] Mojito Tie VideoCrafter2 [5] Mojito Tie OpenSora plan [23]
@@
Table 3: Quantitative evaluation of different designs for the Motion Intensity Modulator (MIM). Combining motion intensity embeddings with text embeddings achieves the best overall performance, balancing video quality, semantic similarity, and motion alignment. Treating motion intensity embeddings as a global conditional input shows good motion alignment but significantly degrades video quality
@@
Qualitative Comparison on Directional Control
We present a qualitative comparison with Tora [50] in Figure 3. The text prompt used is “A red helium balloon floats slowly upward into the sky over a barren desert and expansive Gobi landscape.” The benefits of Mojito are threefold. First, Mojito features a training-free capability, enabling precise directional control at test time without requiring additional training data or fine-tuning. Unlike training-based approaches like Tora, Mojito can dynamically adjust object trajectories during sampling at inference time, ensuring robust control without added computational cost in training. Second, Mojito provides fine-grained directional control by allowing users to specify motion paths for named objects. This level of control contrasts with baseline models like Tora, which lack the capability to follow object-specific directional prompts and typically follow arbitrary trajectories. In Figure 3, we showcase scenarios where Mojito is given specific object names (e.g., “red helium balloon”) and precise locations, enabling it to guide each object’s motion along user-defined paths.
Human Evaluation.
We conduct human evaluations to assess the effectiveness of Mojito in both motion intensity control and directional control. For evaluating the Motion Intensity Modulator, we compare Mojito with baselines by rewriting the text prompts for baselines to explicitly specify the desired intensity level. The results of this evaluation are presented in Table 1. Details on the evaluation and settings are provided in the Supplementary Materials.
@@ 5.5 Visualization of Attention Maps
To understand the progression of directional control in Mojito, we examine cross-attention maps at different sampling steps within the denoising network. The text prompt is “A small red ball bounces along a cobblestone path lined with greenery.” As shown in Figure 4, the attention maps for the target object (the “red ball””) initially display random patterns at the beginning of the sampling process. However, as sampling steps progress, the attention gradually refines and starts to align with the bounding boxes defined for each video timestep. This refinement ensures that the generated video closely follows the user-defined trajectory.
@@ 5.6 Ablation Study
5.6.1 Directional Control
In this section, we analyze the effects of various configurations for directional control, specifically examining different values of guidance strength ($\eta$), temporal smoothness loss ($\lambda$), and guidance steps. The results are presented in Figure 5. The text prompt used is “A vintage wooden sailboat glides steadily down a mist-covered river.” and “A dolphin leaps through the waves of the ocean at sunrise.”
Temporal Smoothness Loss.
When $\lambda=0$ (no temporal smoothness loss), the generated video exhibits noticeable changes in views and inconsistent sailboat appearances.
Guidance Strength.
With $\eta=0$ (no guidance strength), the generated “sailboat” does not follow specified bounding boxes at all. As $\eta$ increases, the sailboat begins to follow bounding boxes more closely. The best results reaches with $\eta=100$, where the sailboat’s trajectory aligns well with the bounding boxes while maintaining good visual quality. However, setting $\eta$ to an excessively high value (e.g., $\eta=10000$) results in severe degradation of video quality, with visible artifacts such as square patterns.
Guidance Steps.
We perform an additional experiment to evaluate the effect of varying the number of guidance steps on video quality and trajectory alignment. The results, shown in Table 2, indicate that increasing the number of guidance steps from $t=1$ to $t=10$ generally enhances trajectory alignment without significantly impacting video quality. However, when the number of guidance steps becomes too large (e.g., $t\geq 30$), the video quality deteriorates substantially. Qualitative results illustrating this effect are provided in Figure 6.
5.6.2 Motion Intensity Modulator
@@
Without MIM.
Removing the MIM module during training and testing. In this setup, we perform inference by rewriting the text prompts to include the intensity level.
Global Conditional Input.
Combining the intensity embedding with FPS and time embeddings, using this global embedding as a conditioning input for convolution layers.
Direct Text Input with Fine-tuning.
In this approach, we rewrite the text prompts to include the motion intensity level and then fine-tune the model using this new data. Both training and inference are guided by the modified text input that explicitly specifies the motion intensity.
@@ 6 Conclusion
In this work, we presented Mojito, a novel framework for text-to-video generation that enables precise control over motion direction and intensity. By introducing the Directional Motion Control module and the Motion Intensity Modulator module,  Mojito allows to specify object trajectories and motion intensity, enhancing flexibility in video generation. Experiments demonstrate Mojito’s effectiveness and efficiency in producing high-quality, motion-controlled videos with minimal reliance on additional annotated videos.
Supplementary Material
@@
7 Model Architecture
We illustrate the architecture of Mojito in Figure 7. The model builds upon a VAE-based latent diffusion framework and incorporates both spatial and temporal transformers to achieve fine-grained control over motion direction and intensity. Spatial transformers process spatial information within each frame, while temporal transformers ensure temporal coherence across video frames.
8 Dataset
We use Panda 70M [8], InternVID [42], and Mira [21] datasets to train Mojito for motion intensity control. For InternVID and Mira, we follow the official data processing setups.
For Panda 70M, we process the data in two stages to ensure semantic consistency and quality. In the first stage, we split long videos using PySceneDetect with a content detector and a cut-scene threshold of 25 frames, introducing artificial cuts every 5 seconds for long, uncut footage. To filter clips with semantic transitions, we compute ImageBind [12] features at 10% and 90% of each clip and retain only those with a Euclidean distance $\leq 1.0$.
9 Implementation Details
9.1 Training Details
We train our model using a transformer-based diffusion framework. The training process incorporates a combination of spatial and temporal transformers to ensure both spatial coherence and motion control across video frames. Key training configurations are detailed below:
9.1.1 Training Configuration
The model uses a learning rate of $1\times 10^{-3}$ without scaling and is optimized with a step-based LambdaLR scheduler. The scheduler initiates a linear decay of the learning rate from the first training step, targeting a final decay ratio of 0.01 over 20,000 steps. Gradient clipping is applied with a maximum norm of 0.5. Mixed precision training is employed to improve efficiency, with DeepSpeed stage-2 optimization enabled for memory management.
9.1.2 Model Architecture
The autoencoder processes input frames at a resolution of $320\times 512$, downsampling them into a latent space with a spatial resolution of $40\times 64$ and a channel depth of 4. The conditional input is encoded using a FrozenOpenCLIPEmbedder [18], which remains fixed during training to leverage text prompts effectively [20]. The model also integrates additional conditioning inputs, such as FPS information, to improve motion control.
9.1.3 Motion Conditioning and Loss Function
Motion intensity control is achieved by introducing optical flow-based motion embeddings as additional conditions. The motion embeddings are combined with text embeddings and fed into cross-attention layers. The diffusion process optimizes a noise-prediction objective:
$\mathcal{L}=\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\|_{2}^{2},$ (8)
where $\epsilon_{\theta}$ predicts the added noise, and $c$ represents the combined text and motion intensity conditions.
9.1.4 Training Strategy
The model is trained for a maximum of 5,000,000 steps with a batch size of 1 and gradient accumulation over 2 steps. Logging occurs every 50 steps, and checkpoints are saved every 1,000 steps, with additional periodic saving every 5,000 steps. Callback functions log intermediate outputs to monitor the training process.
We utilize PyTorch Lightning for distributed training with DeepSpeed Stage-2 optimization. The training is benchmarked for efficiency, using 32-bit precision for stability. Memory optimizations include overlapping communication and computation, ensuring efficient usage of GPU resources.
@@ 9.1.4 Training Strategy
10 Human Evaluation Interface

### v2

@@ 4.4 Model Backbone
The backbone of Mojito is based on a Conv-Spatial-Temporal Transformer architecture. The model structure consists of input blocks, a mid-block, and output blocks, each containing convolutional layers, temporal transformers, and spatial transformers. This design enables handling motion control in both the temporal and spatial dimensions. The details of model backbone are given in the Appendix.
The Temporal Transformer processes video data along the time axis, capturing dependencies between frames by reshaping the input into temporal sequences. This block includes relative position encoding and optional causal masking for unidirectional attention, making it effective for capturing motion patterns across frames; The Spatial Transformer captures spatial relationships within each frame. Using linear or convolutional projections, this block reshapes and processes spatial information to learn localized patterns, which is essential for high-resolution video generation. Cross-attention layers enable alignment with external context when available, further refining spatial coherence in the generated content. The Directional Motion Control (DMC) module operates within Spatial Transformer to enforce trajectory-based motion guidance.
@@ 5.1 Datasets
We use Panda70M (Chen et al., 2024c), InternVID (Wang et al., 2023c), and Mira (Ju et al., 2024) to train both the Mojito backbone and the Motion Intensity Modulator module. From Panda-70M, we extract a high-quality 10M subset for training. Additionally, we process 1,000 videos from Panda-70M for evaluation of motion intensity modulation. For motion direction guidance evaluation, we utilize the LaSOT dataset (Fan et al., 2021), which provides paired text descriptions, bounding box annotations, and corresponding video sequences. For InternVID and Mira, we process and use the entire dataset for training. Further details on dataset preprocessing and usage can be found in the Appendix.
Figure 4: Qualitative comparison of directional control with Tora.  Mojito achieves motion control comparable to Tora while offering additional capabilities to specify objects and precise locations without training. The red bounding boxes, serving as inputs to Mojito, guide the balloon to follow the specified trajectory.
@@ 5.2 Evaluation Metrics
For quantitative evaluation, we employ commonly used metrics: Fréchet Video Distance (FVD) (Unterthiner et al., 2019) and CLIPSIM (Xing et al., 2025), which calculates the average similarity across all generated frames in relation to the input caption. Specifically, we compute the FVD for a set of 300 videos, each consisting of 16 frames.
To evaluate the effectiveness of directional guidance, we utilize the off-the-shelf OWL-ViT-large open-vocabulary object detector (Minderer et al., 2022) to extract bounding box annotations from the generated videos. We then compute the Intersection-over-Union (IoU) (Rezatofighi et al., 2019) between the detected and target bounding boxes to measure spatial alignment. We also compute the Centroid Distance (CD) as the distance between the centroid of generated objects and input bboxes, normalized to 1. For motion intensity control, we compute the difference between the detected average optical flow in the generated video and the target motion intensity.
@@ 5.3 Implementation Details
Model Backbone. For the Variational Autoencoder (VAE) framework, the encoder processes input frames at a resolution of 320$\times$512, progressively downsampling through a series of residual blocks and convolutional layers with channel multipliers of $(1,2,4,4)$. This results in a latent representation with a spatial resolution of $40\times 64$. We use CLIP (ViT-H-14) (Radford et al., 2021) as the text encoder, allowing for high-quality text-to-video alignment. For video processing, spatial and temporal downsampling factors of 8 are applied in both dimensions. The attention matrices (such as $q$ and $kv$) within the spatial transformer are structured as $(t,b,c)$, where $t$ represents the temporal length. Trajectory adjustments are applied only within the mid-block and the initial layers of the upsampling branch in the denoising U-Net (Ronneberger et al., 2015), which provides a balance between maintaining video quality and achieving precise motion control. Additional implementation details and parameter settings are provided in the Appendix.
Training. We use the DeepSpeed (Rasley et al., 2020) framework to train our model, utilizing memory and computation optimizations for distributed training. The model training is carried out using the FusedLamb (You et al., 2019) optimizer combined with the LambdaLRScheduler. We use a batch size of 4 and a base learning rate of $1.0\times 10^{-5}$. More details can be found in the Appendix.
@@ 5.4 Performance Evaluation
Table 1: Human evaluation between Mojito, OpenSora (Zheng et al., 2024), OpenSora plan (Lab & etc., 2024), and VideoCrafter2 (Chen et al., 2024a). It includes 2400 comparisons on 400 video pairs. Each video pair contains 5 ratings from different human annotators on Amazon Turk. The detailed setup is given in the Supplementary Materials.
Mojito Tie OpenSora Mojito Tie VideoCrafter2 Mojito Tie OpenSora plan
@@
Table 3: Quantitative evaluation of different designs for MIM. Combining motion intensity embeddings with text embeddings achieves the best overall performance, balancing video quality, semantic similarity, and motion alignment. Treating motion intensity embeddings as a global conditional input shows good motion alignment but significantly degrades video quality.
@@
Qualitative Evaluations on Directional Control. We present a qualitative comparison with Tora (Zhang et al., 2024) in Figure 4. The text prompt used is “A red helium balloon floats slowly upward into the sky over a barren desert and expansive Gobi landscape.” The benefits of Mojito are threefold. First, Mojito features a training-free capability, enabling precise directional control at test time without requiring additional training data or fine-tuning. Unlike training-based approaches like Tora, Mojito can dynamically adjust object trajectories during sampling at inference time, ensuring robust control without added computational cost in training. Second, Mojito provides fine-grained directional control by allowing users to specify motion paths for named objects. This level of control contrasts with baseline models like Tora, which lack the capability to follow object-specific directional prompts and typically follow arbitrary trajectories. In Figure 4, we showcase scenarios where Mojito is given specific object names (e.g., “red helium balloon”) and precise locations, enabling it to guide each object’s motion along user-defined paths. Third, the DMC module in Mojito is model-agnostic and easily pluggable into different architectures, including DiT-based models, with the results shown in Figure 7.
Human Evaluation. We conduct human evaluations to assess the effectiveness of Mojito in both motion intensity control and directional control. For evaluating the Motion Intensity Modulator, we compare Mojito with baselines by rewriting the text prompts for baselines to explicitly specify the desired intensity level. The results of this evaluation are presented in Table 1. Details on the evaluation and settings are provided in the Appendix.
@@ 5.5 Visualization of Attention Maps
To understand the progression of directional control in Mojito, we examine cross-attention maps at different sampling steps within the denoising network. The text prompt is “A small red ball bounces along a cobblestone path lined with greenery.” As shown in Figure 8, the attention maps for the target object (the “red ball”) initially display random patterns at the beginning of the sampling process. However, as sampling steps progress, the attention gradually refines and starts to align with the bounding boxes defined for each video frame step. This refinement ensures that the generated video closely follows the user-defined trajectory.
@@ 5.6 Ablation Study
5.6.1 Designs in Directional Motion Control
In this section, we analyze the effects of various configurations for directional control, specifically examining different values of guidance strength ($\eta$), temporal smoothness function ($\lambda$), and guidance steps. The results are presented in Figure 5. The text prompt used is “A vintage wooden sailboat glides steadily down a mist-covered river.” and “A dolphin leaps through the waves of the ocean at sunrise.”
Temporal Smoothness Function. When $\lambda=0$ (no temporal smoothness function), the generated video exhibits noticeable changes in views and inconsistent sailboat appearances.
Guidance Strength. With $\eta=0$ (no guidance strength), the generated “sailboat” does not follow specified bounding boxes at all. As $\eta$ increases, the sailboat begins to follow bounding boxes more closely. The best results reaches with $\eta=100$, where the sailboat’s trajectory aligns well with the bounding boxes while maintaining good visual quality. However, setting $\eta$ to an excessively high value (e.g., $\eta=10000$) results in severe degradation of video quality, with visible artifacts such as square patterns.
Guidance Steps. We perform an additional experiment to evaluate the effect of varying the number of guidance steps on video quality and trajectory alignment. The results, shown in Table 2, indicate that increasing the number of guidance steps from $t=1$ to $t=10$ generally enhances trajectory alignment without significantly impacting video quality. However, when the number of guidance steps becomes too large (e.g., $t\geq 30$), the video quality deteriorates substantially. Qualitative results illustrating this effect are provided in Figure 6.
5.6.2 Alternative Choices for Motion Intensity Modulator
@@
Without MIM. Removing MIM during training and testing. In this setup, we perform inference by rewriting the text prompts to include the intensity level.
Global Conditional Input. Combining the intensity embedding with FPS and time embeddings, using this global embedding as a conditioning input for convolution layers.
Direct Text Input with Fine-tuning. In this approach, we rewrite the text prompts to include the motion intensity level and then fine-tune the model using this new data. Both training and inference are guided by the modified text input that explicitly specifies the motion intensity.
@@ 6 Conclusion
In this work, we presented Mojito, a novel framework for text-to-video generation that enables integrated control over motion trajectory/direction and intensity. By introducing the Directional Motion Control module and the Motion Intensity Modulator,  Mojito allows specifying object trajectories and motion intensity during inference. The DMC module is training-free, while the MIM module does not require additional motion data annotations, making it the first approach to explicitly model motion intensity in text-to-video generation and can be trained efficiently. Experimental results demonstrate that Mojito effectively produces high-quality, motion-controlled videos while remaining generalizable and orthogonal to other existing approaches.
Impact Statement
In this paper, the video generation framework may be misused to generate misleading or harmful content (Mirsky & Lee, 2022) and raises concerns about fairness and potential biases in generated outputs. To mitigate these risks, Mojito can be deployed with safeguards such as watermarking generated content, implementing harmful video filters to ensure appropriate usage, and addressing biases in training data to promote fairness and inclusivity in its applications.
Appendix A Model Architecture
We illustrate the architecture of Mojito backbone in Figure 9. The model builds upon a VAE-based latent diffusion framework and incorporates both spatial and temporal transformers within the U-Net architecture to achieve fine-grained control over motion direction and intensity. Spatial transformers process spatial information within each frame, while temporal transformers ensure temporal coherence across video frames.
Appendix B Additional Qualitative Results with Cosmos
We provide additional examples generated by applying the DMC module on Cosmos (NVIDIA et al., 2025) using the 7B diffusion version in Figure 7. The objects in the generated videos strictly follow the input bounding boxes and trajectory guidance, indicating the DMC module is model-agnostic and easy to plug in.
Figure 7: Additional examples demonstrate the application of the DMC module on the DiT-based video generation backbone, Cosmos (NVIDIA et al., 2025). The model successfully follows both bounding box guidance and trajectory constraints
Appendix C Visualizations on Attention Map
We conduct more visualizations on cross-attention maps at different sampling steps within the denoising network in Figure 8. The attention maps for the target object (the “red ball”) initially display random patterns at the beginning of the sampling process. However, as sampling steps progress, the attention gradually refines and starts to align with the bounding boxes defined for each video frame step.
Appendix D Algorithm Framework
The detailed algorithm framework for the DMC module is presented in Algo. 1.
Appendix E Dataset
We use Panda 70M (Chen et al., 2024c), InternVID (Wang et al., 2023c), and Mira (Ju et al., 2024) datasets to train Mojito for motion intensity control. For InternVID and Mira, we follow the official data processing setups.
For Panda 70M, we process the data in two stages to ensure semantic consistency and quality. In the first stage, we split long videos using PySceneDetect with a content detector and a cut-scene threshold of 25 frames, introducing artificial cuts every 5 seconds for long, uncut footage. To filter clips with semantic transitions, we compute ImageBind (Girdhar et al., 2023) features at 10% and 90% of each clip and retain only those with a Euclidean distance $\leq 1.0$.
Appendix F Implementation Details
F.1 Training Details
For the Mojito backbone, we utilize the pretrained VAE encoder from (Chen et al., 2024a). The training process integrates spatial and temporal transformers alongside the MIM module to ensure spatial consistency and controlled motion across video frames.
The model is trained for a maximum of 5,000,000 steps with a batch size of 1 and gradient accumulation over 2 steps. Logging occurs every 50 steps, and checkpoints are saved every 1,000 steps, with additional periodic saving every 5,000 steps to evaluate and pick the best model checkpoint.
We utilize PyTorch Lightning for distributed training with DeepSpeed Stage-2 optimization, using 32-bit precision for stability. Memory optimizations include overlapping communication and computation, ensuring efficient usage of GPU resources. We adopt a fixed learning rate of $1\times 10^{-3}$, optimized using a step-based LambdaLR scheduler. The scheduler applies a linear decay from the initial step, reducing the learning rate to 1% of its original value over 20,000 steps. Gradient clipping is enforced with a maximum norm of 0.5 to stabilize training. Mixed precision training is employed for computational efficiency, with DeepSpeed stage-2 optimization enabled for effective memory management. The model is trained on a large-scale cluster using 16 nodes with a total of 128 A100 GPUs.
F.2 Model Architecture Configuration
The autoencoder processes input frames at a resolution of $320\times 512$, downsampling them into a latent space with a spatial resolution of $40\times 64$ and a channel depth of 4. The conditional input is encoded using a FrozenOpenCLIPEmbedder (Ilharco et al., 2021), which remains fixed during training to leverage text prompts effectively for generation. The model also integrates additional conditioning inputs, such as FPS information, to improve motion control.
Input : Input text $y$, object phrase $y_{n}$, user-defined trajectory $\mathcal{T}$, guidance strength $\eta$, temporal smoothness weight $\lambda$, final guidance timestep $t_{g}$
Output : Generated video from $\boldsymbol{z}^{(0)}$
1
2Initialize latent variable $\boldsymbol{z}^{(T)}\sim\mathcal{N}(0,I)$ for $t\leftarrow T$ to $t_{g}$ do
3       Compute noise level $\sigma_{t}=\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}$
4      Compute attention scores: $A_{i}^{(\gamma)}=\text{Attention}(y_{n},u,i)$
5      Define trajectory-aligned region: $B_{i}=\text{Region}(\mathcal{T},i)$
      Compute energy function: $E_{i}\left(A_{i}^{(\gamma)},B_{i},n\right)=\left(1-\frac{\sum_{u\in B_{i}}A_{i,u,n}^{(\gamma)}}{\sum_{u}A_{i,u,n}^{(\gamma)}}\right)^{2}$  ;
        $\triangleright$ Eq. 4
6
      Compute temporal smoothness loss: $\mathcal{T}_{s}=\mathbb{E}_{i}\left[\|A_{i,:,:}^{(\gamma)}-A_{i-1,:,:}^{(\gamma)}\|^{2}\right]$  ;
        $\triangleright$ Eq. 5
7
8      Compute total gradient: $G=\nabla_{\boldsymbol{z}^{(t)}}\sum_{i=1}^{L}\sum_{\gamma\in\Gamma}\left(E_{i}\left(A_{i}^{(\gamma)},B_{i},n\right)+\lambda\mathcal{T}_{s}\right)$
9      Update latent variable: $\boldsymbol{z}^{(t)}\leftarrow\boldsymbol{z}^{(t)}-\sigma_{t}^{2}\eta G$
10 end for
11
/* Perform directional guidance steps until $t_{g}$/
return Generated video from $\boldsymbol{z}^{(0)}$
Algorithm 1 Training-Free Directional Motion Control
@@ 9      Update latent variable: $\boldsymbol{z}^{(t)}\leftarrow\boldsymbol{z}^{(t)}-\sigma_{t}^{2}\eta G$
Appendix G Human Evaluation Interface